if a woman gave birth at 21 how old woud she and the baby be when she is three times older than the child
Answer by ptaylor(2198)

Let x=time that elapses until the woman is 3 times older than the child
So, our equation to solve is:
3x=21+x subtract x from each side
2x=21
x=10.5 years
CK
3*10.5=21+10.5
31.5=31.5
Hope this helps----ptaylor